Comfi Airways, Inc., a small two-plane passenger airline, has asked for your assistance in some basic analysis of its operations. Both planes seat 10 passengers each, and they fly commuters from Comfi’s base airport to the major city in the state, Metropolis. Each month, 40 round-trip flights are made. Shown below is a recent month’s activity in the form of a cost-volume-profit income statement.

Fare revenues (400 passenger flights) $48,000
Variable costs
    Fuel $17,210
    Snacks and drinks 760
    Landing fees 2,100
    Supplies and forms 1,050 21,120
Contribution margin 26,880
Fixed costs
    Depreciation 2,900
    Salaries 15,260
    Advertising 400
    Airport hanger fees 1,600 20,160
Net income $6,720

a) Calculate break even point in dollars

b)Calculate the break even point in number of passenger flights

c) Without calculations, determine the contribution margin at the break even point

c2) If ticket prices were decreased y 10%, passenger flights would increase by 25%. However total variable costs would increase by the same percentage as passenger flights

How much would net income be impacted by this change?

Net income__________increase/decrease to $_____________

Should the ticket price decrease be adopted?
Answer #1
Break even point = Fixed cost/contribution margin
Contribution margin =    Contribution margin
Sales
26,880 = 56%
48,000
1) Break even point 20160/56%
$ 36000 Answer
2) Break even 300 Flights
Fixed cost/contribution margin per unit
contribution margin per unit= contribution/passenger flights
26880/400
67.2
3) Break even point 20,160 answer
(At break even point contribution is equal to fixed cost
4) Ticket prices 48000/400= 120
now decrease in price 10%
New ticket price 108
passenger flights 400*125%= 500
New variable cost 52.8*125%= 66
New contribution margin 108-66
42
total contribution margin (500*42) 21000
less fixed cost 20,160
net income(new) 840
net income will decrease by 5880 (original net income - new net income)